The footnpag Package

Joachim Schrod^{∗}

August 4, 1995 (Revision 2.5 of footnpag.sty)

1. This package is appropriate for numbering footnotes separately on each page. It may be used with all standard document classes (and I assume with all other well written ones). ‘Numbering’ here does not only mean supplying Arabic numbers, if your class or another package sets up appropriate symbols, they are used instead.

You just have to use the footnpag package, the rest will happen automagically. You will need two LaTEX runs, as with cross references or citations.

2. Known Problems.

Unlike with references, LaTEX will not issue a warning if a footnote number is incorrect due to new page breaks. Thus run LaTEX always another time if you want a final document.

This is scheduled to change in the next revision.

3. An auxiliary file is used, named jobname.fot. This is not necessary by itself, it would have been better to use the aux file. (Actually, this package was written for plain TEX originally, there is no aux file there.)

This is also scheduled to change in the next revision. In fact, it’s the same problem like the previous one: When I use the aux file, LaTEX’s warning mechanism can be utilized.

4. There is no possibility to change the number that’s used for the first footnote on a page. Some people like to start their footnotes with ‘0’, there exists even a package for that

named zero. footnpag should cooperate with that package.

If one uses symbols as footnote markers, the first footnote is marked with an asterisk. One might want to use a dagger instead, that’s not possible currently.

Most probably this feature will be made available via the new “keyword-value” option scheme, as you’ll find it in the graphicx package.

∗Email: schrod@iti.informatik.th-darmstadt.de

Revision Log for footnpag-user.tex

Revision 2.1 (created at August 4, 1995 by Joachim Schrod)

Made a LaTEX 2ε package from this style option.

User manual is a separate document now, that’s better for installation. Started to change the distribution into one that conforms to the ‘supported bundle guidelines.’